Status Report

NASA Space Station On-Orbit Status 31 Jan 2004

By SpaceRef Editor
February 2, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 31 Jan 2004

All ISS systems continue to function nominally, except as noted previously or below   

Progress 13P (M1-11 or No. 260) docking to the Service Module (SM) aft port was successfully accomplished on time at 8:13am EST on automatic control.   [The entire process of fully automated rendezvous, closure, final approach and capture, followed by closing of hooks and latches, went smoothly and without issues.  The Progress approached from forward, starboard and nadir for the period from 1 km into the beginning of flyaround.  The Lab window shutter had to remain closed when the Progress was within 915 m of ISS, which occurred at ~7:44am.]

After hard-dock, pressurization of the Progress-to-SM vestibule and installation of the quick-disconnect screw clamps (with SM thrusters disabled), the crew conducted a one-hour leak check of the interface vestibule, accompanied by air sampling in the SM for ammonia with the IPD Draeger tubes sampler through the Progress vestibule hatch valve (and also later during the actual hatch opening).

CDR/SO Michael Foale reconfigured the communications setup, deactivated the Ku-band TV system used for covering the approach and docking, and demated the UOP (utility outlet panel) bypass power cable from the RWS DCP (robotics workstation display & control panel).

Before the docking, Alex Kaleri worked on the SM’s Kurs-P hardware, connecting it to the AKR antenna-3.

Opening of the two hatches is scheduled for ~12:40pm, supported by a teleconference with the ground for Go from MCC-M/TsUP.   [The crew has special safety instructions regarding Progress ingress and ESA’s HEAT payload, which contains toxic ammonia (NH3).  Removal of HEAT is scheduled for 2/3 (Tuesday).  If ammonia is released in the ISS, the crew will retreat to a safe haven, potentially overnight, while the TCCS (trace contaminant control subsystem) will scrub the ammonia to acceptable levels.  The Russian  BMP (harmful impurities removal system) will remove the small amounts of NH3 that may be in the safe haven, and the crew will use Drager tubes to verify that NH3 levels are at a low enough concentration to doff masks.]

After hatch opening, FE Alexander Kaleri will use the Russian AK-1M instrument to perform the obligatory air sampling in the Progress, then install the threaded QD (quick disconnect) screw clamps of the SM’s docking and internal transfer mechanism (SSVP), which rigidize the mating surfaces.  Sasha will then deactivate the cargo ship and install the ventilation air duct in the transfer tunnel.  Tomorrow (Sunday), he is to remove the probe-and-cone docking mechanism to clear the passage for the ensuing transfer operations, starting tomorrow.

Structural vibrations during the docking were recorded with the SDMS (structural dynamic measurement system).  SAMS (space acceleration measurement system) and MAMS (microgravity acceleration measurement system) also took data.

The CDR deactivated the EarthKAM experiment.  The hardware stow is planned for 2/2 (Monday).   [EarthKAM reports that the activity was very successful, involving 57 schools, with over 700 images taken.]

The crew worked out on TVIS, RED exerciser and VELO cycle with load trainer.

Mike Foale conducted the daily routine maintenance of the SM’s SOZh life support system.

Mike also completed the regular task of transferring data files from the physical exercise equipment to the MEC (medical equipment computer) via memory card and RED log entries, for downlink on OCA comm.  Later, he performed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleting them on the HRM.

The CDR performed the regular once-a-week maintenance reboot on the operational PCS laptops and also restarted the OCA comm router laptop (every two weeks).

During the 13P rendezvous phase the crew performed a test of the TORU teleoperated approach and docking system, as called for by Russian procedure. The TORU test was performed twice but failed both times to establish a radio link between the SM and the Progress.  Prior to docking, once Progress came into the comm window over RGS (Russian ground sites), TsUP commanded the Progress to use the second string (backup) TORU transceiver.  This resulted in a functional TORU link between SM and Progress.

MCC-M will be investigating the cause of the TORU test failures.

Weekly Science Update (Expedition Eight — 13th):
GASMAP:   The next scheduled health check should occur in early February.

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound:   Planned.

Hand Posture Analyzer (HPA:  Looking forward to future operations.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   PFMI is back in business and looking forward to continued operations.
Space Acceleration Measurement System (SAMS):   Mike was thanked for rebooting the SAMS ICU this week in time to capture data for the 12P undocking.
Microgravity Acceleration Measurement System (MAMS):   MAMS captured 12P undocking and 13P docking.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):  PromISS OBT and part one of the PromISS setup are complete.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Renal Stone (RS):   The crew’s last in-flight data collection session will occur in early April.  Thanks to the crew for their continued support with taking their daily pills at dinner time.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   Pre-Increment requirements have been completed.  A second BBT (Beacon & Beacon Tester) session will be scheduled in US Lab.

Foot/Ground Reaction Forces during Space Flight (FOOT):   Data downlink is tentatively planned for mid-February.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Looking forward to the next set of FDI Tissue Culture runs.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Fifty-one schools, including two from Japan and one from Chile, are ready for EarthKAM ops next week, to be conducted from the SM.

Earth Science Toward Exploration Research (ESTER):  Looking forward to the sessions next year.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   Thank you to Mike for powering up of CGBA this week.  So far all parameters are nominal.  The ground looks forward to the stowing of GCF (Granada Crystallization Facility) into CGBA tomorrow morning, and to the start of the Yeast GAP experiment next week.

Miscible Fluids in Microgravity (MFMG):    Nothing new.

Educational Payload Operations (EPO):  The education demonstrations continue to be excellent.  The video will be used in a variety of ways.  The ground is looking forward to upcoming activities.

Crew Earth Observations (CEO):   One of the recent CEO images of the large Upsala Glacier of the Southern Patgonian Ice Field has been selected for publication in Earth Observatory.  The article highlights visible changes noted along the glacier’s terminus when compared with views taken 4 years ago by the crew of Increment 1.   Expedition 8 has continued to take excellent images of both the two large icebergs northeast of South Georgia Island as well as very useful views of the several ongoing plankton blooms in the South Atlantic.  The crew’s technique of including land features as reference points in the images greatly increase their potential value.  A survey of their most recently downlinked images through January 28, indicates that they have been making heavy use of both the 400mm lens and the doubler.   The ground has noticed that these latest images are slightly soft.  The 400 with the doubler is a challenging combination to get good focus with.  Low light and atmospheric turbidity will also degrade image quality.  The ground recommended that the crew concentrate on high contrast targets for best results.

Today’s CEO (crew earth observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Tropical Cyclone Frank (Dynamic event. Heading southwest towards Mauritius, Frank is strengthening and is expected to be a category 4 storm by Sunday.  ISS passed west of the eye, now visible on satellite imagery), Beijing, China (looking right for the city center), Delhi, India (Nadir pass over this city [13.8 million in 2001], located on the Yamuna River, the main headwater river of the Ganges.  The city can be difficult to detect: looking for a point on the river where the crew could see Delhi’s radiating highway transport network), Addis Ababa, Ethiopia (there are no thunderstorms on the Ethiopian plateau.  Looking right of track for this capital city.  The ground knows of no good images of Addis), Plankton, Patagonia (looking both sides of track for ~2.5 mins., especially north of the Falkland Islands), and Cairo, Egypt (nadir pass over this city of ~12.5 million).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

ISS Orbit  (as of this morning, 7:50am EST [= epoch]):

  • Mean altitude — 368.4 km
  • Apogee — 373.7 km
  • Perigee — 363.1 km
  • Period — 91.9 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0007838
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 90 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29672

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.